A simple pendulum is found to vibrate 50 times within 200 s. When 1.5 m of its length is reduced to a certain length, it vibrates 50 times in 175 s. Find the original length of the pendulum. 


1
Expert's answer
2022-02-02T05:15:15-0500

Given:

"N_1=50,\; t_1=200\: \rm s"

"N_2=50,\; t_2=175\: \rm s"

"l_1=l"

"l_2=l-1.5"


The period of oscillation of a simple pendulum is given by

"T=\frac{t}{N}=2\pi\sqrt{l/g}"

Hence

"\frac{t_1}{N_1}=2\pi\sqrt{l/g}"

"\frac{t_2}{N_2}=2\pi\sqrt{(l-1.5)/g}"

We get

"(t_1\/t_2)^2=l\/(l-1.5)"

"64\/49=l\/(l-1.5)"

"l=6.4\:\rm m"